OpenStudy (anonymous):

I'LL GIVE METAL! Contestants in a game show are each given three pairs of numbers (for example, 2-3, 5-6, and 6-1). A person in the audience is randomly selected to roll two dice, one white and one red, where white symbolizes the first number in the pair and red the second number. What is the probability that each player has of winning?

OpenStudy (ledah):

are these the normal 6 side dice? or are we talking something a little different? just thought i'd ask that first

OpenStudy (anonymous):

Normal dice

OpenStudy (anonymous):

I'm actaully not 100% sure the question is odd..

OpenStudy (ledah):

yeah that is a strange question. but lets assume they mean a normal set of dice.

OpenStudy (anonymous):

I think it might be 1/2

OpenStudy (ledah):

ok, thats a start, what makes you think that its 1/2?

OpenStudy (anonymous):

the probability of rolling an even number is 1/2.

OpenStudy (ledah):

so an even number? not every number? well with even numbers and dice you actually have 3 even numbers which are, 2,4,6 and with a set of dice it doubles your chances.

OpenStudy (anonymous):

the options are 1/2, 1/3, 2/3, 1/12, 1/5

OpenStudy (ledah):

Alright give me one second. ^3^ I'll get right back on this

OpenStudy (anonymous):

Okay, I appriciate it. :)

OpenStudy (ledah):

alright so, can you give me an explanation as to why you think that answer is right, cause i even questioned myself on this one.

OpenStudy (ledah):

I can't even give a proper explanantion for this. but no matter what. your answer is Correct. P(A)= n(A)/n(S)= 18/36= 1/2 =50%

OpenStudy (ledah):

Look you answered your own question o_o that's awesome lol

OpenStudy (anonymous):

haha thank you! odd question I know.

OpenStudy (ledah):

well you got it. so ^^ no worries. You did great for going into something so odd